According to the school librarian, the average number of pages of books in the reference section is 240. To

test her claim, she collected a sample of 15 books and after noting the number of pages of each book, she

determined that the mean number of pages is 224.6 with a standard deviation of 4.1. At α = 0.01, will the

librarian be able to prove her claim?


1
Expert's answer
2022-05-25T12:46:06-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=240"

"H_1:\mu\not=240"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=14" and the critical value for a two-tailed test is "t_c =2.976842."

The rejection region for this two-tailed test is "R = \{t:|t|>2.976842\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{224.6-240}{4.1/\sqrt{15}}=-14.5473"

Since it is observed that "|t|=14.5473>2.976842=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=14" degrees of freedom, "t=-14.5473" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 240, at the "\alpha = 0.01" significance level.